a) If you have Cobb Douglas utility function ……………………………….. ( 7 marks)
U: U (x. y) = x0.5 y0.5
Assume that the consumer is able to spend $40 on goods b, and c. where Px = 1, and Py= 2. Find the maximum quantities of b, and c the consumer can afford to buy and give him the highest utility.
1
Expert's answer
2020-04-14T10:00:42-0400

Solution:


"MU_x=\frac{0.5y^{0.5}}{x^{0.5}}"


"MU_y=\frac{0.5x^{0.5}}{y^{0.5}}"

"p_xq_x+p_yq_y=40"


"x+2y=40"


"\frac{MU_x}{p_x} = \frac {MU_y}{p_y}"


"\frac{0.5y^{0,5}}{x^{0.5}} = \frac{0,5x^{0,5}}{2y^{0,5}}"


"2y=x"

"2y+2y=40"


"y=10"


"x=20"